FLUORESCENCE EXCITATION SPECTRA OF PHOTO-FRAGMENTED NITROBENZENE USING A PI-COSECOND LASER: POTENTIAL EVIDENCE FOR NO PRODUCED BY TWO DISTINCT CHANNELS.

\underline{CHRISTOPHER J. LUE}, CHAKREE TANJAROON, J. BRUCE JOHNSON, SCOTT W. REEVE, Arkansas Center for Laser Applications and Science and Department of Chemistry and Physics, P.O. Box 419 State University, AR 72467; SUSAN D. ALLEN, Embry Riddle Aeronautical University, 600 S. Clyde Morris Boulevard, Daytona Beach, FL 32114.

Upon absorption of a UV photon, nitrobenzene can dissociate into C_{6}H_{5}O and NO through two different mechanisms. Evidence for these mechanisms was obtained from velocity map imaging (VMI) studies and theoretical calculations.^{a} VMI experiments showed NO produced with two distinct rotational distributions, which the calculations explained as a fast and a slow channel for NO production. We have recorded high resolution fluorescence excitation spectra of the NO resulting from photo-fragmented nitrobenzene using a pulsed picosecond tunable laser (pulse width ≈ 15 ps) by means of a two-color process. In the two-color process, photons of a particular energy dissociated the nitrobenzene while photons of a different energy probed the A^{2}Σ^{+} ← X^{2}Π_{(1/2,3/2)} NO band system between 225-260 nm. This laser system allowed us to vary the delay between the photolysis and excitation pulses. At longer delays (>1 ns), we observed an increase in the population of NO, which may be evidence that at least two photolysis channels produce NO. We present the spectra we recorded at various photolysis/probe delays ranging from 0.025 to 1.5 ns. The spectral subtraction method we used to observe the production increase is introduced.

aHause, M. L.; Herath, N.; Zhu, R.; Lin, M. C. and Suits, A. G. Nat Chem, Nature Publishing Group, 2011, 3, 932-937